Journalist Report – February 20th

Sol 8 – Exploring the 8th continent

“The men of Earth came to Mars. They came because they were afraid or unafraid, because they were happy or unhappy, because they felt like Pilgrims or did not feel like Pilgrims. There was a reason for each man.”

– – Chapter 8 of The Martian Chronicles by Ray Bradbury

The members of today’s EVA crew definitely felt like explorers, if not pilgrims, unafraid to take on an unexplored Martian area! We were very excited this morning to launch the photogrammetry experiment by sending out the first of six EVA crews. Alice, the EVA leader, followed by Quentin and Adrien, headed out to North Ridge to begin 3D-mapping a specific area for other EVA crews to visit later during the mission. The objective is to compare how performance during an EVA (for example, how fast an astronaut can find a designated spot in unknown Martian terrain) is affected by showing the crew members a 3D map of the area beforehand. After today’s “reconnaissance EVA”, two other crews will go to the same location: one will have been given a 3D-map generated thanks to drone imagery, and another will only be given a 2D map. After arriving on-site, the crew members had to immerse themselves in the space to determine the best place from which to launch our Parrot drone. It was then flown over the chosen area by Quentin while Alice decided which spots would be the target areas for the next crew.

Since I will be part of the next EVA returning to the same location, I was therefore not allowed to see the photogrammetry render for myself, but was told it turned out amazing! Quentin was surprised to see that the generated map extended far wider than expected, beyond the area actually explored by the astronauts. The landscape is well-defined, with all rocks more than a meter high detected and rendered with great precision.

This EVA’s other objectives were also completed: Alice gathered some samples for her geology experiment, and the crew ran some more tests on MegaAres and retrieved data from the weather station before heading to North Ridge. In sum, apart from a few communication issues, the EVA was successful!

After debriefing the EVA during lunch, Alexandre reviewed the schedule to find time to solve a problem with MegaAres: after today’s tests, it was confirmed that part of the metallic sphere is not connected to the rest of the electrical circuit; tomorrow morning’s EVA will be necessary to retrieve the spherical part of the antenna to “debug it”.

It felt good to find ourselves all working together in the Upper Deck of the Hab as the afternoon progressed; I feel like we are becoming more and more accustomed to life as a crew, and considering each other as team members and co-workers, but also as friends.

EVA Report – February 20th

Crew 275 EVA Report 20Feb2023

EVA # 6

Author: Alice Chapiron, EVA Leader and Crew Scientist

Purpose of EVA: The purpose of the EVA was to do photogrammetry of North Ridge with drones.

Start time: 9:02

End time: 11:50


09:02: Airlock closed.

09:09: Rovers Curiosity and Spirit unplugged. We then depart from the hab at 09:12.

09:14: Rovers parked at the right location near the Mega-Ares, METEO and Purple Air instruments.

09:30: Weather Station and Purple Air data copied. Batteries changed. New recording started. Mega-Ares data copied.

09:36: Beginning of the Mega-Ares test. We didn’t have time to finish all the test, so we preferred to continue with the EVA as it was not the main purpose of today.

09:50: Departure from the Mega-Ares, METEO and Purple Air instrument.

09:58: We had to stop to conduct radio tests because the EVA Leader had her antenna inside her helmet.

10:06: Parked at the right location, near the beginning of the gateway to Candor. Walking towards the desired location.

10:10: Arrival on site for the photogrammetry experiments.

10:23: Launching the drone and starting to place the markers coupled with the sample collection.

10:37: The hat of the Crew Scientist fell off, which slightly impaired her vision. The crew botanist secured her way to a safe place and then the problem was solved by removing the hat from the helmet. It was solved in 5 minutes.

11:01: The crew Scientist microphone got off her head. She could not hear clearly the others, neither talk to them. The Crew Botanist took the head of communication. They gathered with the Crew engineer who was nearby.

11:10: The crew Scientist put back the microphone to solve the problem, so the communication could continue safely. We went back to our activities.

11:20: Loss of contact. Both the EVA Leader and the Crew Botanist couldn’t speak nor be heard. The problem was certainly coming from the fact we were all on channel 1 following the previous problem. This was solved after running some tests and being advised by both HabCom and Mission Support. Departure from North Ridge towards the rover.

11:30: Back to the Rover

11:41: Arrival at the hab

11:50: End of the EVA

We forgot one small item on the trunk of Opportunity. It will stay there for the night and be retrieved tomorrow.

Note: At around 11:00, both the Crew Engineer and the Crew Botanist saw 4 pick-ups with around 10 people. They approached around 150 m from us and took pictures. We were really impacted and stressed by this event, while we had to remain focused on our tasks at the same time.

Destination: North Ridge (first we stop where the MegaAres and the METEO station are located)

Coordinates: N4252000 E519000

Participants: EVA Leader: Alice Chapiron (Crew Scientist), Quentin Royer (Crew Engineer), and Adrien Tison (Crew Botanist)

Road(s) and routes per MDRS Map: Cow Dung Road until North Ridge, near the intersection with “Gateway to Candor”, we then walked towards North-West for an estimated 600 meters from the rover.

(See map on the attached picture)

Mode of travel: Driving using Curiosity and Spirit

GreenHab Report – February 20th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 80.5 °F

Average humidity: 17 % (to be double-checked)

Hours of supplemental light: 4 hours

Daily water usage for crops: 10.84 gallons

Daily water usage for research and/or other purposes: 0.26 gallons

Water in the blue tank: 97.13 gallons

Time(s) of watering for crops: 9:53 & 13:57

Changes to crops: N/A

Narrative: This morning I was on an EVA so it was our crew journalist that maintained and watered plants and crops in the morning.

As I came back, I went to check on it at 13:57. Everything is in place and doing well.

I began our Microgreens experiment by taking two pots and planting new microgreens to monitor their growth.

Our environmental sensors have just been deployed so I will be able to double-check the humidity tomorrow.

Apart from that, nothing special to signal, everything looks fine!

Harvest: N/A

Support/supplies needed: N/A

Sol Summary Report – February 20th

Crew 275 Sol Summary Report 20Feb2023

Sol: 8

Summary Title: Sol 8 – Exploring the eighth continent

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today we conducted the first EVA dedicated to photogrammetry. It started very successfully with a stop at the location of the atmospheric instruments to retrieve some data, change batteries, and conduct supplementary tests on Mega-Ares. The EVA crew then reached North Ridge to start their activities: sample collection and 3D mapping of the terrain using drones. If the scientific objectives were completely fulfilled, it is important to make a few remarks regarding this EVA. 1) One crewmember had a problem with their hat and headset during the operations. There was a misunderstanding regarding the actions to take in this situation, which led to a few minutes before taking the decision to put the hand inside the helmet to solve the issue. This event was discussed as a group and the procedure is now clear for everyone. 2) There were issues regarding radio communications within the crew and between HabCom and the Crew. This led to some confusion. Again, this event was debriefed and the procedure is also clear for everyone now. It includes some reminders on how to communicate efficiently on the different radio channels. 3) The EVA crew was approached by 4 pick-ups getting as close as 150 m from them. They took pictures of the crew conducting their activities. This was reported as a very stressful event by the EVA crew, as this took a lot of their mental space as they were piloting a drone at this moment.

Look Ahead Plan:

The tests conducted during the last two EVAs have shown that there was a small issue with the antenna of Mega-Ares, which can be solved only from within the Hab. For this reason, we would like to perform an EVA tomorrow (Sol 9) morning to retrieve this antenna and proceed to a few additional checks on the atmospheric instruments. We plan to put this antenna back in its place on Wednesday (Sol 10) afternoon and to take this opportunity to change the batteries of all the atmospheric instruments.

With the success of the first photogrammetry EVA, a 3D map of the target area has been rendered. Now, we would like to send two EVA crews to the exact same location to check the added value of the 3D map to ease the ability of astronauts to identify some specific features onsite. We thus plan to send an EVA crew to North Ridge with a 3D map on Wednesday morning (Sol 10) and another one with only a 2D map on Thursday morning (Sol 11).

Anomalies in work: None.

Weather: Sunny, low -8°C/18°F, high 6°C/42°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Operations Report – February 20th

Crew 275 Operations Report 20Feb2023
SOL: 8

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet

Notes on non-nominal systems: None

Spirit rover used: Yes
• Hours: 209.1
• Beginning charge: 100%
• Ending charge: 66%
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.1
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 221.0
• Beginning charge: 100%
• Ending charge: 68%
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.5
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report. The EVA Crew forgot a camera lens protector when they came back, on the edge of the trunk of Opportunity rover. Mission Support was advised by radio at the end of the EVA.

Summary of Hab operations: The Crew conducted routine operations. The Lower Deck heater was ON all day. If the external temperature remains the same as today, we will switch it off as of tomorrow.
• WATER USE: 10.30 gallons
• Water: 355.35 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: The EVA Crew experienced an issue with the radio at North Ridge. After issues of communication between the HabCom and the EVA Crew, everyone in the EVA Crew was on Channel 1, and they were unable to communicate between each other for about one minute. Back on channel 1 (EVA leader) and 2 (the 2 other EVA Crew Members), the EVA Crew was able to communicate again. We have no explanation for this, and we will conduct further tests.

Summary of GreenHab operations: The Crew botanist conducted routine operations and planted two microgreens crops to start his experiment. An environmental sensor was placed in the GreenHab.
• WATER USE: 11.10 gallons
• Heater: On
• Supplemental light: From 10:00 pm to 2:00 am
• Harvest: No

Summary of ScienceDome operations: The Crew Scientist conducted routine operations. An environmental sensor was placed in the ScienceDome.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today. An environmental sensor was placed in the RAM.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: A Crew Member slipped in the stairs, and slightly hurt their elbow. The HSO handled the situation and there is nothing to worry about.

Questions, concerns, and requests to Mission Support: None.

Copyright © The Mars Society. All rights reserved. | Main Site